# Topology and Algebra of Bonded Knots and Braids

### v2

@@ 8. The theory of enhanced bonded links & braids
Figure 75. Interpretation of enhanced bonds.
@@ 8.2. The enhanced bonded braid group
Recall that bonded braids are defined as classical braids equipped with embedded horizontal simple arcs, the bonds. Similarly, enhanced bonded braids are classical braids equipped with two different types of bonds, the attracting and the repelling bonds. By abuse of notation, we shall denote by $b_{i}$ an attracting elementary enhanced bond and by $b_{i}^{-1}$ we shall denote the corresponding repelling bond. Moreover, we will denote by $b_{i,j}$ an attracting bond between the $i^{th}$ and $j^{th}$ strands of an enhanced bonded braid and by $b_{i,j}^{-1}$ the corresponding repelling bond, such that when we have two consecutive bonds of different types between the same strands of an enhanced bonded braid, they cancel out (see Figure 79).
@@
The above results lead to the definition of the enhanced bonded braid group $EB_{n}$ with operation the concatenation of bonded braids in the set of enhanced bonded braids on $n$-strand. Note that as in the case of bonded braids, we will denote by $b_{i,j}$ all possible sequences of $j-i+1$ o’s and u’s, that indicate the type of the crossings formed between the enhanced bond and the strands of the braid, and by $b_{i,j}^{-1}$ its corresponding inverse in the enhanced bonded braid group. Finally, isotopy between two enhanced bonded braids is defined in the same way as isotopy between bonded braids, since these isotopy moves do not alter the type on the enhanced bonds. Then with the same reasoning as earlier we obtain:
@@
The enhanced tight bonded braid group $EB_{n}$ is the group generated by $\sigma_{1},\ldots,\sigma_{n-1}$, the attracting bonds $b_{1},\ldots,b_{n-1}$ and the repelling bonds $b_{1}^{-1},\ldots,b_{n-1}^{-1}$, collectively called enhanced bonds (see Figure 78), with operation the usual braid concatenation. The generators satisfy the same relations as the generators of the tight bonded braid monoid, $BB_{n}$, together with the following relations:
[…]
@@
The enhanced tight bonded braid group $EB_{n}$ admits an irredundant presentation with generators the classical braid generators $\sigma_{1},\ldots,\sigma_{n-1}$ together with a single attracting bond generator $b_{1}$ and its inverse $b_{1}^{-1}$, subject to the relations:
$\begin{array}[]{rcll}\sigma_{i}\,\sigma_{j}&=&\sigma_{j}\,\sigma_{i}&\quad\text{for }|i-j|>1,\\[4.0pt] \sigma_{i}\,\sigma_{i+1}\,\sigma_{i}&=&\sigma_{i+1}\,\sigma_{i}\,\sigma_{i+1}&\quad\text{for all }i,\\[4.0pt] b_{1}\,\sigma_{j}&=&\sigma_{j}\,b_{1}&\quad\text{for }j>2,\\[4.0pt] b_{1}\,(\sigma_{2}\sigma_{1})(\sigma_{1}\,\sigma_{2})&=&(\sigma_{2}\sigma_{1})(\sigma_{1}\,\sigma_{2})\,b_{1}&\\[4.0pt] b_{1}\,\sigma_{1}&=&\sigma_{1}\,b_{1}.&\end{array}$
8.3. The analogues of Alexander and Markov theorems for enhanced bonded braids
@@
Before moving toward a braid equivalence theorem for enhanced bonded braids, we first need to generalize the notion of the $L$-moves on this setting. We define an $L$-move on an enhanced bonded braid as in Definition 7.1, by ignoring the types of the bonds (recall Figure 64). It follows from the results § 7 that with the use of the $L$-moves on enhanced bonded braids, we obtain the $L$-move equivalence for enhanced bonded braids, by treating enhanced bonds as simple bonds. Namely, we obtain the following:
Theorem 8.7 ($L$-equivalence for enhanced bonded braids).
Two enhanced bonded braids have isotopic closures if and only if one can be obtained from the other by a finite sequence of the following moves together with enhanced bonded braid isotopy:
$\begin{array}[]{lllcll}1.&L-{moves}&&&&\\[4.0pt] 2.&{Enhanced\ Bond\ Conjugation:}&\alpha&\sim&b_{i}^{\pm 1}\,\alpha\,b_{i}^{\mp 1},&{for}\ \alpha,\,b_{i}\in EB_{n}.\\ \end{array}$
We replace moves (1) of Theorem 7.3 by considering stabilization moves and conjugation for classical braids and we obtain the analogue of the Markov theorem for enhanced bonded braids. Namely:
Theorem 8.8 (Markov equivalence for enhanced bonded braids).
Two enhanced bonded braids have isotopic closures if and only if one can be obtained from the other by a finite sequence of the following moves together with enhanced bonded braid isotopy:
$\begin{array}[]{lllcll}1.&{Conjugation:}&\alpha&\sim&\beta^{\pm 1}\,\alpha\,\beta^{\mp 1},&{for}\ \alpha,\,\beta\in EB_{n},\\[4.0pt] 2.&{Markov\ Stabilization:}&\alpha&\sim&\alpha\,\sigma_{n}^{\pm 1},&{for}\ \alpha\in EB_{n}.\\ \end{array}$
@@ 9.1. Bonded knotoids
Knotoids were introduced by Turaev in [41] as a generalization of 1-1 tangles by allowing the endpoints to be in different regions of the diagram. Namely: a knotoid diagram $K$ in an oriented surface $\Sigma$ is a generic immersion of the unit interval [0,1][0,1][ 0 , 1 ] into $\Sigma$ whose only singularities are transversal double points endowed with over/undercrossing data called crossings. The images of 0 and 111 under this immersion are called the endpoints of $K$ (leg and head of $K$ respectively) and are distinct from each other and from the double points (see Figure 81 by ignoring the bonds).
A knotoid in $\Sigma$ is then an equivalence class of knotoid diagrams in $\Sigma$ up to the equivalence relation induced by the Reidemeister moves (recall Figures 58), that take place away from the endpoints of the knotoid. A multi-knotoid diagram in $\Sigma$ is a generic immersion of the unit interval [0,1][0,1][ 0 , 1 ] and a number of copies of $S^{1}$ into $\Sigma$. Note that in the theory of multi-knotoids, the pulling of a strand that is adjacent to an endpoint is not allowed, since this would result into trivial knotoid diagrams. These moves are illustrated in Figure 80, and they are called forbidden moves of multi-knotoids. A knotoid has a natural orientation from leg to head. A multi-knotoid diagram is oriented if orientations are also assigned to its closed components.
@@
[…]
@@ 9.2. Types of closures for bonded multi-knotoids
[…]
@@
We first treat strands of the braids that contain the nodes of bonds as shown in Figures 62 and 63. In that way we obtain up-arcs that do not contain the nodes of bonds. We apply then the braidoiding algorithm of [18] (which is similar to the braidoiding algorithm we presented in § 6 and we obtain a bonded braidoid, whose closure is isotopic to the bonded multi-knotoid. ∎
10. Further Work
@@
We will formulate $L$-equivalence in the setting of bonded braidoids, extending the correspondences to the bonded and enhanced bonded braidoid categories. We will define the plat closure for bonded braids and braidoids and will formulate Hilden and Birman type theorems [21, 6].
We will explore a bonded Morse category for multi-knotoids and linkoids. Bonds are included just as with braids, but in the Morse category maxima (caps) and minima (cups) are allowed in the diagrams. Thus there are generators for a braided tensor category consisting in local bonds, crossings of two types, cups and caps and identity lines. There is an extension of Reidemeister moves for this Morse category. Then one has representation theorems that assure that entities can be written in Morse form, and generalized Reidemeister moves imply that topologically equivalent entities can be connected by sequences of moves. This categorical approach allows the construction of invariants of bonded Morse diagrams in terms of functors from the Morse category to a category of modules and traces. For a formulation of a Morse category for knotoids see [17]. We will detail this approach in a subsequent paper.

### v4

@@ 8. The theory of enhanced bonded links & braids
Figure 76. Interpretation of enhanced bonds as ‘attracting’ and ‘repelling’.
@@ 8.2. The enhanced bonded braid group
Recall that bonded braids are defined as classical braids equipped with embedded horizontal simple arcs, the bonds. Similarly, enhanced bonded braids are classical braids equipped with two different types of bonds, the attracting and the repelling bonds. By abuse of notation, we shall denote by $b_{i}$ an attracting elementary enhanced bond and by $b_{i}^{-1}$ we shall denote the corresponding repelling bond. Moreover, we will denote by $b_{i,j}$ an attracting bond between the $i^{th}$ and $j^{th}$ strands of an enhanced bonded braid and by $b_{i,j}^{-1}$ the corresponding repelling bond, such that when we have two consecutive bonds of different types between the same strands of an enhanced bonded braid, they cancel out (see Figures 77 and 80).
@@
The above results lead to the definition of the enhanced bonded braid group $EB_{n}$ with operation the concatenation of bonded braids in the set of enhanced bonded braids on $n$-strand. Note that as in the case of bonded braids, we will denote by $b_{i,j}$ all possible sequences of ji+1j-i+1 o’s and u’s, that indicate the type of the crossings formed between the enhanced bond and the strands of the braid, and by $b_{i,j}^{-1}$ its corresponding inverse in the enhanced bonded braid group. Finally, isotopy between two enhanced bonded braids is defined in the same way as isotopy between bonded braids, since these isotopy moves do not alter the type on the enhanced bonds. Then with the same reasoning as earlier we obtain:
@@
The enhanced tight bonded braid group $EB_{n}$ is the group generated by the classical invertible braid generators $\sigma_{1},\ldots,\sigma_{n-1}$, the attracting bonds $b_{1},\ldots,b_{n-1}$ and the repelling bonds $b_{1}^{-1},\ldots,b_{n-1}^{-1}$, collectively called enhanced bonds (see Figure 79), with operation the usual braid concatenation. The generators satisfy the same relations as the generators of the tight bonded braid monoid, $BB_{n}$, together with the following relations:
$b_{i}\,b_{i}^{-1}\ =\ 1\ =\ \ b_{i}^{-1}\,b_{i},\ {\rm for\ all}\ i.$
@@
The enhanced tight bonded braid group $EB_{n}$ admits an irredundant presentation with generators the classical invertible braid generators $\sigma_{1},\ldots,\sigma_{n-1}$ together with a single attracting bond generator $b_{1}$ and its inverse $b_{1}^{-1}$, subject to the relations:
[…]
8.3. Analogues of Alexander and Markov theorems for enhanced bonded braids
@@
Before moving toward a braid equivalence theorem for enhanced bonded braids, we first need to generalize the notion of the $L$-moves on this setting. We define an $L$-move on an enhanced bonded braid as in Definitions 7.1 and 7.4, by ignoring the types of the bonds (recall Figure 65). It follows from the results § 7 that with the use of the $L$-moves on enhanced bonded braids, we obtain the $L$-move equivalence for enhanced bonded braids, by treating enhanced bonds as simple bonds. Namely, we obtain the following:
Theorem 8.7 (Bonded $L$-equivalence for enhanced bonded braids).
Two enhanced bonded braids upon closure give rise to topologically vertex isotopic oriented enhanced standard bonded links if and only if they can be obtained one from the other by a finite sequence of enhanced bonded braid isotopy and the following moves:
[…]
Furthermore, two enhanced tight bonded braids upon closure give rise to topologically vertex isotopic oriented enhanced tight bonded links if and only if they can be obtained one from the other by a finite sequence of tight bonded braid isotopy and the following moves:
[…]
We can replace moves (1) of Theorem 8.7 by considering stabilization moves and conjugation for classical braids and, along with the algebraization of the bonded $L$-moves, we obtain the analogue of the Markov theorem for enhanced bonded braids. This is the subject of a sequel work.
@@ 9.1. Bonded knotoids
Knotoids were introduced by Turaev in [38] as a generalization of 1-1 tangles by allowing the endpoints to be in different regions of the diagram. Namely: a knotoid diagram $K$ in an oriented surface $\Sigma$ is a generic immersion of the unit interval $[0,1]$ into $\Sigma$ whose only singularities are transversal double points endowed with over/undercrossing data called crossings. The images of 0 and $1$ under this immersion are called the endpoints of $K$ (leg and head of $K$ respectively) and are distinct from each other and from the double points (see Figure 82 by ignoring the bonds).
A knotoid in $\Sigma$ is then an equivalence class of knotoid diagrams in $\Sigma$ up to the equivalence relation induced by the Reidemeister moves (recall Figures 58), that take place away from the endpoints of the knotoid. A multi-knotoid diagram in $\Sigma$ is a generic immersion of the unit interval $[0,1]$ and a number of copies of $S^{1}$ into $\Sigma$. Note that in the theory of multi-knotoids, the pulling of a strand that is adjacent to an endpoint is not allowed, since this would result into trivial knotoid diagrams. These moves are illustrated in Figure 81, and they are called forbidden moves of multi-knotoids. A knotoid has a natural orientation from leg to head. A multi-knotoid diagram is oriented if orientations are also assigned to its closed components.
@@
A bonded multi-knotoid diagram in $\Sigma$ is a pair $(K,B)$ of a knotoid $K$ equipped with a set $B$ of bonds, defined as in the case of bonded links in Definition 1.1. View Figure 82. Further, a bonded linkoid diagram is defined to be an immersion of a disjoint union of finitely many unit intervals whose images are knotoid diagrams, equipped with bonds.
@@ 9.2. Types of closures for bonded multi-knotoids
A braidoid diagram $B$ is a system of a finite number of arcs immersed in $0,1]\times[0,1]\subset\mathbb{R}^{2}$, where $\mathbb{R}^{2}$ is identified with the xt-plane, such that the t-axis is directed downward. The arcs of $B$ are the strands of $B$. Each strand is naturally oriented downward, with no local maxima or minima, following the natural orientation of $[0,1]$. Moreover, there are only finitely many intersection points among the strands, which are transversal double points endowed with over/under data, and are called crossings of $B$. A braidoid diagram has two types of strands, the classical strands, i.e. braid strands connecting points on $0,1]\times\{0\}$ to points on $0,1]\times\{1\}$, and the free strands that either connect a point in $0,1]\times\{0\}$ or in $0,1]\times\{1\}$ to an endpoint located anywhere in $0,1]\times[0,1]$, or they connect two endpoints that are located anywhere in $0,1]\times[0,1]$. These points that lie on $0,1]\times\{0\}$ or $0,1]\times\{1\}$, are called braidoid ends.
@@
We first treat strands of the braids that contain the nodes of bonds as shown in Figures 63 and 64. In that way we obtain up-arcs that do not contain the nodes of bonds. We apply then the braidoiding algorithm of [17] (which is similar to the braidoiding algorithm we presented in § 6 and we obtain a bonded braidoid, whose closure is isotopic to the bonded multi-knotoid. ∎
10. Conclusions and Further Work
In this paper we have used topological, diagrammatic and algebraic methods to define and analyze bonded structures. This means that it is assumed that the reader understands how to translate such language into three dimensions. Diagrams can be regarded as schemata for producing specific three dimensional structures and embeddings, just as a weaving pattern is an instruction for producing a given weave. Our topological argumentation is fully rigorous because the diagrammatics are a formal system for these topological and combinatorial structures. Thus the diagrammatics form a pivotal place where one has formality on the one hand and interpretability on the other.
This topologists’ stance may be new to some scientists who look directly at three dimensional structure. Our approach is particularly useful for the formulation of algebraic invariants and for the formulation of algorithms for computation of invariants. To understand the full story for subjects like protein folding, one needs three dimensional structure and here the diagrams form a basis for further articulation. If we wish to further study physical interactions (beyond the present paper), then one can add more structure to the combinatorial models given here and work with them three dimensionally. This is a project for further research, and it promises deeper relationships between our invariants and the physical behaviours of molecules.
The combinatorial and algebraic coding of structures that we use in this paper has many potentials for applications. The standard diagramming can be used to produce embeddings in three dimensional space directly. The braid representations are concise algebraic methods to encode bonded structures and they are new and need to be studied further for their potential. Each method of formalizing a three dimensional topological structure has its own properties that deserve further research.
In a paper to follow the present work we will investigate more deeply the invariants described here. These include the unplugging invariants, the invariants using insertion and the Kauffman bracket and Jones polynomials, and the new invariants that arise from the $L$-move braiding formulations for bonded knots and links. In the case of the $L$-moves, we will continue the formulation in terms of generalizations of the algebraic Markov Theorem for bonded braids and consequent invariants defined in these terms.
Computational approaches to our work include the use of molecular databases (we refer to Sulkowska’s works in KnotProt and the Protein Data Base [8]) where one can translate three dimensional data into the combinatorics, allowing one to compute invariants and analyse the bonding structure of proteins. In the other direction one can translate diagrammatic encoding into chosen three dimensional embeddings and then work with these models in three dimensional space, including comparison with molecules from the database. This relationship with experimental and three dimensional information is ongoing. We have used these methods in our previous work and will continue to use it in applications of the present research.
In the case of the ideas we have suggested for Feynman diagrams, the ideas are in a state of flux in that it is not yet clear that adding knotted structure to Feynman diagrams will advance the understanding of quantum field theory. On the other hand, field theoretic approaches to protein folding are clearly needed (and under investigation [35]) and we need to see how such approaches are related to the three dimensional combinatorial and topological structure of molecules. The Feynman diagrams are an intermediation between the field theory and the combinatorics. For all these reasons we regard the present work as constructing a foundation for much-needed further research.
Bonded knotoids are especially relevant for modeling open chains such as proteins. We will revisit (enhanced) bonded knotoids and braidoids and their topological interactions in the form of (enhanced) bonded braidoid equivalences.
@@ 10. Conclusions and Further Work
In another direction we will define the plat closure for bonded braids and braidoids and will formulate Hilden and Birman type theorems for turning an unoriented bonded knot/knotoid to bonded plat/platoid and for their equivalences.
We will explore a bonded Morse category, a natural next step for multi-knotoids and linkoids: a Morse-theoretic diagrammatics that admits cups/caps and incorporates bonds. This framework will carry a corresponding move calculus and normal-form results, enabling functorial invariants via monoidal functors to module categories (with trace constructions).